An object is to be dropped from an aeroplane so as to hit a target on the ground. If the aeroplane is travelling horizontally at 80m/s at a height of 2000m. Calculate
(a)the position of the aeroplane when the object is released.
(b)the time for the object to hit the target from the instant of release.
1
Expert's answer
2020-08-31T12:47:33-0400

a)


"x=vt=(80)(20)=1600\ m"

b)


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(2000)}{9.8}}=20\ s"